OpenStudy (anonymous):

Y over -2 plus 5 equal 13

OpenStudy (lsdxmdmaddicthc):

Set it up in an equation : (Y/-2) + 5 = 13

OpenStudy (lsdxmdmaddicthc):

Now you want to get Y alone on one side of the equation

OpenStudy (lsdxmdmaddicthc):

Do you know how to do that?

OpenStudy (anonymous):

14 + 12y ≤ 50 12y ≤ 36 y ≤ 3

OpenStudy (anonymous):

no

OpenStudy (lsdxmdmaddicthc):

I can explain it if you want?

OpenStudy (anonymous):

yes please

OpenStudy (lsdxmdmaddicthc):

Okay. So you have your equation set up : (Y/-2) + 5 = 13

OpenStudy (lsdxmdmaddicthc):

Whatever you do to one side you must do the other side.

OpenStudy (lsdxmdmaddicthc):

So if we want to get Y alone, we would have to remove the rest of that side...

OpenStudy (anonymous):

yes....y+10=13 idk

OpenStudy (lsdxmdmaddicthc):

To remove the +5, we would subtract 5 to cancel it out. Since we are subtracting 5 from one side, we must also subtract it from the other side.

OpenStudy (lsdxmdmaddicthc):

So it would look like this : (Y/-2) + 5 = 13 (Y/-2) + 5 - 5 = 13 - 5 (Y/-2) = 8

OpenStudy (lsdxmdmaddicthc):

Do you understand it so far?

OpenStudy (anonymous):

y/-2=8

OpenStudy (anonymous):

yes

OpenStudy (lsdxmdmaddicthc):

Yes, so now we want to get rid of the /-2

OpenStudy (lsdxmdmaddicthc):

So we would multiply by -2. It would look like this Y/-2 * -2 = 8 * -2

OpenStudy (lsdxmdmaddicthc):

I think you can figure out the value of Y from here. I'll tell you if you're correct.

OpenStudy (anonymous):

thanks for helping me understand it

OpenStudy (anonymous):

-16

OpenStudy (lsdxmdmaddicthc):

You're welcome! Math can seem really confusing, I know. But it is actually pretty interesting once you get someone to explain it :)

OpenStudy (lsdxmdmaddicthc):

Yep, that's correct. Glad I could help!